# Improving the solver for the Balitsky-Kovchegov evolution equation with Automatic Differentiation

### v1

@@ Abstract
The Balitsky-Kovchegov (BK) evolution equation is an equation derived from perturbative Quantum Chromodynamics that allows one to calculate the scattering amplitude of a pair of quark and antiquark off a hadron target, called the dipole amplitude, as a function of the collision energy. The initial condition, being a non-perturbative object, usually has to be modeled separately. Typically, the model contains several tunable parameters that are determined by fitting to experimental data. In this contribution, we propose an implementation of the BK solver using differentiable programming. Automatic differentiation offers the possibility that the first and second derivatives of the amplitude with respect to the initial condition parameters are automatically calculated at all stages of the simulation. This fact should considerably facilitate and speed up the fitting step. Moreover, in the context of Transverse Momentum Distributions (TMD), we demonstrate that automatic differentiation can be used to obtain the first and second derivatives of the amplitude with respect to the quark-antiquark separation. These derivatives can be used to relate various TMD functions to the dipole amplitude. Our C++ code for the solver, which is available in a public repository [1], includes the Balitsky one-loop running coupling prescription and the kinematic constraint. This version of the BK equation is widely used in the small-$x$ evolution framework.
@@ 1 Introduction
It is expected that new measurements of cross sections in $e+\textrm{A}$ collisions conducted at the future Electron-Ion Collider [2, 3] in the low-$x$ regime will shed more light on the long-sought phenomenon of gluon saturation. It would manifest itself by the fact that the rise of the cross section with energy, predicted by the linear evolution equation derived with perturbative QCD [4, 5], is replaced by some sort of flattening due to the non-linear effects, i.e. gluon recombination, captured by the Balitsky-Kovchegov (BK) equation [6, 7, 8, 9] or more generally the Jalilian-Marian-Iancu-McLerran-Weigert-Leonidov-Kovner (JIMWLK) equation [10, 11, 12, 13, 14, 15, 16]. The calculation of the Deep Inelastic Scattering (DIS) cross section requires a non-perturbative model of the initial configuration of color sources in the target that is then evolved to higher energies with one of the above equations. Typically one uses the Golec-Biernat–Wusthoff (GBW) model [17, 18] or McLerran-Venugopalan (MV) model [19, 20, 21] to obtain the initial distribution of color charges. In either case, the values of the free, non-perturbative parameters of these models must be determined by comparison with experimental data. In the most naive approach one constructs an appropriate $\chi^{2}$ function that is then minimized numerically with respect to the model parameters. Such a numerical minimization often requires the evaluation of gradients in the parameter space of the $\chi^{2}$ function in order to approach the minimum. The derivatives can only be accessed if the estimates of $\chi^{2}$ are available for at least two different values of the initial parameters, i.e. for each parameter an additional solution of the evolution equation is needed. This procedure must be performed at each step of the iterative evolution.
@@
Although there exist automatic tools that, given a simulation software, generate a new version where the calculation of derivatives is automatically implemented [22], we provide our own implementation. This is motivated, first of all, by the fact that we are interested in the first derivatives and the Hessian matrix of only a few specific parameters which may appear independently in different primitives of the program. Our second motivation is to automatically estimate the first and second derivatives with respect to the dipole size $\bm{\bar{r}}_{\perp}$ of the dipole amplitude. However, the BK equation is an integro-differential equation for the dipole amplitude $S_{\bm{\bar{r}}_{\perp}=\bm{\bar{x}}_{\perp}-\bm{\bar{y}}_{\perp}}(\eta)$ which contains a convolution with respect to the daughter dipoles’ sizes $\bm{r}_{xz}$ and $\bm{r}_{zy}$ where $\bm{\bar{r}}_{\perp}=\bm{r}_{xz}+\bm{r}_{zy}$. Thus, if we choose one daughter dipole size $\bm{r}_{xz}$ as the independent integration variable, the dependence of the other daughter dipole size $\bm{r}_{zy}=\bm{r}_{zy}(\bm{\bar{r}}_{\perp})$ becomes non-trivial. Therefore, it is not straightforward and automatic to track down the differentiable variable $\bm{\bar{r}}_{\perp}$ in the various places of the source code and we cannot rely on automatic tools.
In the following we provide more details about the leading-order (LO) BK equation and the more advanced form that includes several phenomenologically relevant improvements that we used in our work. We also provide a brief discussion of the convergence of the BK equation and its derivatives with respect to dipole size. In Section 3, we discuss the changes necessary to implement Automatic Differentiation into the framework that solves the BK equation. In Section 4, we show the comparison of our results against the existing software. We take this opportunity to provide some remarks on the performance. Section 5 is devoted to the applications. First, we show the quality of the derivatives of the DIS cross section with respect to the initial condition parameters. Second, we present functions $\mathcal{K}_{+}$ and $\mathcal{K}_{-}$ needed for the calculation of the TMD distributions in the Gaussian approximation. These functions are constructed from derivatives with respect to the quark-antiquark separation of the dipole amplitude.By comparing the results obtained with automatic differentiation against the divided differences method applied directly to the dipole amplitude, we provide evidence for the superiority of the AD approach. Eventually, we give conclusions and outlook in Section 6.
The code described in this work is publicly available in the source code repository [1].
@@ 2 Balitsky-Kovchegov evolution equation
where $\bar{\alpha}_{s}$ is the strong coupling constant of QCD taken here as a constant, $\eta$ is the rapidity variable, and $S_{\bm{\bar{x}}_{\perp}\bm{\bar{y}}_{\perp}}(\eta)$ is the dipole amplitude describing the amplitude of scattering the pair of quark and anti-quark of size $\bm{\bar{r}}_{\perp}=\bm{\bar{x}}_{\perp}-\bm{\bar{y}}_{\perp}$ on a proton in the limit of very high energies. The dipole kernel reads
@@
By symmetry arguments the dipole amplitude cannot depend on the orientation of the vector $\bm{\bar{r}}_{\perp}$ but only on its length $r=|\bm{\bar{r}}_{\perp}|$, the most convenient way to estimate the integral over $\bm{\bar{z}}_{\perp}$ is to introduce radial coordinates, $r_{z}=|\bm{\bar{x}}_{\perp}-\bm{\bar{z}}_{\perp}|$ and $\phi=\angle(\bm{\bar{z}}_{\perp},\bm{\bar{x}}_{\perp})$. From translational and rotational symmetries and from the assumption of an infinite target it follows that the origin of the coordinate system can be set at $\bm{\bar{x}}_{\perp}$ without any loss of generality. Hence, rewritten in radial variables Eq. (1) reads
[…]
@@
$\frac{\partial S(r,\eta)}{\partial\eta}=\int d\phi\,dr_{z}\,r_{z}\,\left[\frac{\bar{\alpha}_{s}(r)}{2\pi r^{2}_{z}}\left(\frac{r^{2}}{r^{2}_{zy}}+\frac{\bar{\alpha}_{s}(r_{z})}{\bar{\alpha}_{s}(r_{zy})}-1+\frac{r^{2}_{z}}{r^{2}_{zy}}\left(\frac{\bar{\alpha}_{s}(r_{zy})}{\bar{\alpha}_{s}(r_{z})}-1\right)\right)\right]\times\\ \times\big{[}S(r_{z},\eta-\delta_{r_{z};r})S(r_{zy},\eta-\delta_{r_{zy};r})-S(r,\eta)\big{]},$ (4)
@@
$C$ is a tunable parameter and has been set to $3.8078$ [30], while $b_{0}$ is the zeroth coefficient of the $\beta$-function in QCD, $b_{0}=(11N_{c}-2N_{f})/(12\pi)\approx 0.716$. Although $r$ and $r_{z}$ are positive, it may happen that $r_{zy}$ is zero and therefore we introduced $\epsilon_{\alpha_{s}}$ in Eq. (5) as a regularization and we set it to $10^{-7}$.
@@ 2.1 Numeric solution of the BK equation
The typical approach to solve numerically Eq. (1) or Eq. (4) is to discretize the rapidity variable $\eta$ and introduce a discrete grid in the transverse plane for the $\bm{\bar{z}}_{\perp}$ vector. In radial coordinates this amounts to discretizing the radial variable $r_{z}$ and the angle $\phi$. Then, starting with a given analytic model for the initial condition $S(r,\eta=0)$, one uses the Euler scheme to advance the dipole amplitude $S(r,\eta)$ by a small increment $\delta_{\eta}$ in $\eta$. This requires the evaluation of the convolution with respect to $r_{z}$ and $\phi$, where usually $r_{z}$ is discretized on a logarithmic scale.
@@ 2.2 BK equation for the first and second derivative of dipole amplitude
The BK equation is free of ultraviolet divergences, this follows from the probability conservation of the wave function where divergences of virtual diagrams cancel divergences of real diagrams when the emitted gluon $\bm{\bar{z}}_{\perp}$ is at the location of one of the two parents $\bm{\bar{x}}_{\perp}$ or $\bm{\bar{y}}_{\perp}$. Thus, the apparently singular behavior of the dipole kernel Eq. (2) when $\bm{\bar{z}}_{\perp}\rightarrow\bm{\bar{x}}_{\perp}$ or $\bm{\bar{z}}_{\perp}\rightarrow\bm{\bar{y}}_{\perp}$, is compensated by the combination of dipoles distributions $[S_{\bm{\bar{x}}_{\perp}\bm{\bar{z}}_{\perp}}S_{\bm{\bar{z}}_{\perp}\bm{\bar{y}}_{\perp}}-S_{\bm{\bar{x}}_{\perp}\bm{\bar{y}}_{\perp}}]$ in the integrand of Eq. (1). As an illustration, for a Gaussian initial condition of the dipole distribution, the combination of dipoles $[S_{\bm{\bar{x}}_{\perp}\bm{\bar{z}}_{\perp}}S_{\bm{\bar{z}}_{\perp}\bm{\bar{y}}_{\perp}}-S_{\bm{\bar{x}}_{\perp}\bm{\bar{y}}_{\perp}}]$ expanded around small separation $\bm{\bar{z}}_{\perp}-\bm{\bar{x}}_{\perp}$ scales as $\mathcal{O}\left((\bm{\bar{z}}_{\perp}-\bm{\bar{x}}_{\perp})^{2}\right)$, and the integrand in Eq. (1) is free of divergence at said point. This conclusion for UV-finiteness is unchanged when derivatives are taken w.r.t. the parent dipole size $\bm{\bar{r}}_{\perp}$, but due to the Leibniz rule individual terms may appear singular.
In the present case, it is convenient to use the polar coordinates, and the LO dipole kernel would reads:
$\mathcal{M}_{\bm{\bar{x}}_{\perp}\,\bm{\bar{y}}_{\perp}\,\bm{\bar{z}}_{\perp}}=\frac{r^{2}}{r_{z}^{2}r_{zy}^{2}}\longrightarrow\frac{r^{2}}{r_{z}^{2}\,\left(r_{z}^{2}+r^{2}-2rr_{z}\text{cos}(\phi)+\epsilon^{2}\right)}.$ (7)
The regulator $\epsilon=5\cdot 10^{-4}\,\text{GeV}^{-1}$ is introduced to improve the numerical stability of the equations. This is illustrated here for the LO dipole kernel, and implemented for the kernel from [25] in the provided code. We emphasize, that this regulator does not regulate the UV behavior of the equations, since they are already UV finite by themselves. In particular, $\epsilon$ can be smaller than the intrinsic UV scale provided by the lattice spacing. We checked that varying $\epsilon$ in a wide range from $5\cdot 10^{-5}\,\text{GeV}^{-1}$ to $5\cdot 10^{-3}\,\text{GeV}^{-1}$ had negligible impact on the results. On the other hand, $\epsilon$ stabilizes the evaluation of the derivatives of the dipole amplitude w.r.t. $r$ since no $/0$ 0 symbol is encounter in the limit $r_{z}\to r$, $\phi\to 0$. Note that the Jacobian for the polar coordinate provides $r_{z}$ that improves the behavior of the integrand when $r_{z}\rightarrow 0$. Hence, no regulator was needed there to evaluate the dipole amplitude and its first two derivatives w.r.t. $r$.
@@ 3.1 Forward-mode AD and dual numbers
In the present application we employ forward-mode implementation of AD because of its better performance for a small number of input parameters compared to reverse-mode [33, 34]. We replace real numbers with dual numbers defined as a C++ class as shown in Listing 1. All necessary numeric primitives needed for the calculation are rewritten to operate on dual numbers and are contained in a single header file. We use operator overloading in C++ to handle dual numbers in order to increase the readability of the original source code, see Listing 2 for the example of multiplication operator.
//number of parameters with derivatives
const int NN = 2;
//number of independent entries in the hessian matrix
const int NN_hessian = NN*NN - NN*(NN-1)/2;
template <typename T> class dual {
public:
T val;
T derr;
T der2r;
T der[NN];
T der2[NN_hessian];
dual(T v){
val = v;
derr = 0.0;
der2r = 0.0;
for(int i=0; i<NN; i++)
der[i] = 0.0;
for(int i=0; i<NN_hessian; i++)
der2[i] = 0.0;
template <typename T> dual<T> operator*(dual<T> a, dual<T> b){
dual<T> c;
c.val = a.val*b.val;
c.derr = a.derr*b.val + a.val*b.derr;
c.der2r = a.der2r*b.val + 2.0*a.derr*b.derr + a.val*b.der2r;
for(int i=0; i<NN; i++)
c.der[i] = a.der[i]*b.val + a.val*b.der[i];
for(int i=0; i<NN; i++)
for(int j=i; j<NN; j++){
int index = i*NN+j-i*(i+1)/2;
c.der2[index] = a.der2[index]*b.val
+ a.der[i]*b.der[j] + a.der[j]*b.der[i]
+ a.val*b.der2[index];
return c;
It is known from previous studies [35] that the most suitable variables for the evaluation of the convolution in Eq. (1) are polar coordinates together with a logarithmic scale for the radial variable. We implement the Simpson’s 3/8 rule for both the angle and the radial variable. Since we require differentiability with respect to the $\bm{\bar{r}}_{\perp}=\bm{\bar{x}}_{\perp}-\bm{\bar{y}}_{\perp}$ we initialize it using a dual variable as shown in Listing 3.
Listing 3: Declaration of the radial variable. Slogmin is the minimal value of the radius $r_{xy}$ and logstep is the increment on the logarithmic scale. $y$ is the loop iterator.
dual<mytype> r_xy(Slogmin + y*logstep);
r_xy.derr = 1.0/exp(Slogmin + y*logstep);
r_xy.der2r = -1.0/exp(2.0*(Slogmin + y*logstep));
dual<mytype> er_xy = expp(r_xy);

### v2

@@ Abstract
The Balitsky-Kovchegov (BK) evolution equation is an equation derived from perturbative Quantum Chromodynamics that allows one to evolve with collision energy the scattering amplitude of a pair of quark and antiquark off a hadron target, called the dipole amplitude. The initial condition, being a non-perturbative object, usually has to be modeled separately. Typically, the model contains several tunable parameters that are determined by fitting to experimental data. In this contribution, we propose an implementation of the BK solver using differentiable programming. Automatic differentiation offers the possibility that the first and second derivatives of the amplitude with respect to the initial condition parameters are automatically calculated at all stages of the simulation. This fact should considerably facilitate and speed up the fitting step. Moreover, in the context of Transverse Momentum Distributions (TMD), we demonstrate that automatic differentiation can be used to obtain the first and second derivatives of the amplitude with respect to the quark-antiquark separation. These derivatives can be used to relate various TMD functions to the dipole amplitude. Our C++ code for the solver, which is available in a public repository [1], includes the Balitsky one-loop running coupling prescription and the kinematic constraint. This version of the BK equation is widely used in the small-$x$ evolution framework.
@@ 1 Introduction
It is expected that new measurements of cross-sections in $e+\textrm{A}$ collisions conducted at the future Electron-Ion Collider [2, 3] in the low-$x$ regime will shed more light on the long-sought phenomenon of gluon saturation. It would manifest itself by the fact that the rise of the cross-section with energy, predicted by the linear evolution equation derived with perturbative QCD [4, 5], is replaced by some sort of flattening due to the non-linear effects, i.e. gluon recombination, captured by the Balitsky-Kovchegov (BK) equation [6, 7, 8, 9] or more generally the Jalilian-Marian-Iancu-McLerran-Weigert-Leonidov-Kovner (JIMWLK) equation [10, 11, 12, 13, 14, 15, 16]. The calculation of the Deep Inelastic Scattering (DIS) cross section requires a non-perturbative model of the initial configuration of color sources in the target that is then evolved to higher energies with one of the above equations. Typically one uses the Golec-Biernat–Wusthoff (GBW) model [17, 18] or McLerran-Venugopalan (MV) model [19, 20, 21] to obtain the initial distribution of color charges. In either case, the values of the free, non-perturbative parameters of these models must be determined by comparison with experimental data. In the most naive approach, one constructs an appropriate $\chi^{2}$ function that is then minimized numerically with respect to the model parameters. Such a numerical minimization often requires the evaluation of gradients in the parameter space of the $\chi^{2}$ function in order to approach the minimum. The derivatives can only be accessed if the estimates of $\chi^{2}$ are available for at least two different values of the initial parameters, i.e. for each parameter an additional solution of the evolution equation is needed. This procedure must be performed at each step of the iterative evolution.
@@
Although there exist automatic tools that, given a simulation source code, generate a new version where the calculation of derivatives is automatically implemented [22], we provide our own implementation. This is motivated, first of all, by the fact that we are interested in the first derivatives and the Hessian matrix of only a few specific parameters that may appear independently in different primitives of the program. Our second motivation is to automatically estimate the first and second derivatives with respect to the dipole size $\bm{\bar{r}}_{\perp}$ of the dipole amplitude. However, the BK equation is an integrodifferential equation for the dipole amplitude $S_{\bm{\bar{r}}_{\perp}=\bm{\bar{x}}_{\perp}-\bm{\bar{y}}_{\perp}}(\eta)$ which contains a convolution with respect to the daughter dipoles’ sizes $\bm{r}_{xz}$ and $\bm{r}_{zy}$ where $\bm{\bar{r}}_{\perp}=\bm{r}_{xz}+\bm{r}_{zy}$. Thus, if we choose one daughter dipole size $\bm{r}_{xz}$ as the independent integration variable, the dependence of the other daughter dipole size $\bm{r}_{zy}=\bm{r}_{zy}(\bm{\bar{r}}_{\perp})$ becomes non-trivial. Therefore, it is not straightforward and automatic to track down the differentiable variable $\bm{\bar{r}}_{\perp}$ in the various places of the source code and we cannot rely on automatic tools.
In the following, we provide more details about the leading-order (LO) BK equation and the more advanced form that includes several phenomenologically relevant improvements that we used in our work. We also provide a brief discussion of the convergence of the BK equation and its derivatives with respect to dipole size. In Section 3, we discuss the changes necessary to implement Automatic Differentiation into the framework that solves the BK equation. In Section 4, we show the comparison of our results against the existing software. We take this opportunity to provide some remarks on the performance. Section 5 is devoted to the applications. First, we show the quality of the derivatives of the DIS cross-section with respect to the initial condition parameters. Second, we present functions $\mathcal{K}_{+}$ and $\mathcal{K}_{-}$ needed for the calculation of the TMD distributions in the Gaussian approximation. These functions are constructed from derivatives with respect to the quark-antiquark separation of the dipole amplitude. By comparing the results obtained with automatic differentiation against the divided differences method applied directly to the dipole amplitude, we provide evidence for the superiority of the AD approach. Eventually, we give conclusions and outlook in Section 6.
The code described in this work is publicly available in the source code repository [1]: https://bitbucket.org/piotrekkorcyl/bk_with_ad/src/main/.
@@ 2 Balitsky-Kovchegov evolution equation
where $\bar{\alpha}_{s}=\frac{\alpha_{s}N_{c}}{\pi}$, and $\alpha_{s}$ is the strong coupling constant of QCD taken here as a constant, $\eta$ is the rapidity variable, and $S_{\bm{\bar{x}}_{\perp}\bm{\bar{y}}_{\perp}}(\eta)$ is the dipole amplitude describing the amplitude of scattering the pair of quark and anti-quark of size $\bm{\bar{r}}_{\perp}=\bm{\bar{x}}_{\perp}-\bm{\bar{y}}_{\perp}$ on a proton in the limit of very high energies. The dipole kernel reads
@@
In the following, we assume translation and rotational invariance (which also implies the assumption of an infinite target size) [^1] Approaches to BK going beyond assumptions of infinite target size were also developed, see for example [23, 24, 25], but these are beyond scope of this manuscript. so the dipole amplitude cannot depend on the orientation of the vector $\bm{\bar{r}}_{\perp}$ but only on its length $r=|\bm{\bar{r}}_{\perp}|$, $S(r,\eta)$. The most convenient way to estimate the integral over $\bm{\bar{z}}_{\perp}$ is to introduce radial coordinates, $r_{z}=|\bm{\bar{x}}_{\perp}-\bm{\bar{z}}_{\perp}|$ and $\phi=\angle(\bm{\bar{z}}_{\perp},\bm{\bar{x}}_{\perp})$. From assumed symmetries, the origin of the coordinate system can be set at $\bm{\bar{x}}_{\perp}$ without any loss of generality. Hence, rewritten in radial variables Eq. (1) reads
$\frac{\partial S(r,\eta)}{\partial\eta}=\frac{\bar{\alpha}_{s}}{2\pi}\int d\phi\,dr_{z}\,r_{z}\,\frac{r^{2}}{r_{z}^{2}(r^{2}+r_{z}^{2}-2rr_{z}\cos\phi)}\left[S(r_{z},\eta)\,S\left(\sqrt{r^{2}+r_{z}^{2}-2rr_{z}\cos\phi},\eta\right)-S(r,\eta)\right].$ (3)
@@
$\frac{\partial S(r,\eta)}{\partial\eta}=\int d\phi\,dr_{z}\,r_{z}\,\left[\frac{\bar{\alpha}_{s}(r)}{2\pi r^{2}_{z}}\left(\frac{r^{2}}{r^{2}_{zy}+\epsilon^{2}}+\frac{\bar{\alpha}_{s}(r_{z})}{\bar{\alpha}_{s}(r_{zy})}-1+\frac{r^{2}_{z}}{r^{2}_{zy}+\epsilon^{2}}\left(\frac{\bar{\alpha}_{s}(r_{zy})}{\bar{\alpha}_{s}(r_{z})}-1\right)\right)\right]\times\\ \times\big[S(r_{z},\eta-\delta_{r_{z};r})S(r_{zy},\eta-\delta_{r_{zy};r})-S(r,\eta)\big],$ (4)
@@
$C$ is a tunable parameter and has been set to $3.8078$ [31], while $b_{0}$ is the zeroth coefficient of the $\beta$-function in QCD, $b_{0}=(11N_{c}-2N_{f})/(12\pi)\approx 0.716$, where number of colors $N_{c}=3$ and number of quark flavours $N_{f}=3$. Although $r$ and $r_{z}$ are positive, it may happen that $r_{zy}$ is zero and therefore we introduced $\epsilon_{\alpha_{s}}$ in Eq. (5) as a regularization and we set it to $10^{-7}$.
@@ 2.1 Numeric solution of the BK equation
The typical approach to solving numerically Eq. (1) or Eq. (4) is to discretize the rapidity variable $\eta$ and introduce a discrete grid in the transverse plane for the $\bm{\bar{z}}_{\perp}$ vector. In radial coordinates this amounts to discretizing the radial variable $r_{z}$ and the angle $\phi$. It is known from previous studies [32] that the most suitable radial variable for the evaluation of the convolution in Eq. (1) is the logarithmic variable. Then, starting with a given analytic model for the initial condition $S(r,\eta=0)$, one uses the Euler scheme to advance the dipole amplitude $S(r,\eta)$ by a small increment $\delta_{\eta}$ in $\eta$. This requires the evaluation of the convolution with respect to $r_{z}$ and $\phi$, where usually $r_{z}$ is discretized on a logarithmic scale. In our implementation, we used the Simpson $\frac{3}{8}$ rule for both the radial and angular variables. The missing data for the dipole amplitude was obtained by linear interpolation, while the kernel and the running coupling constant were evaluated analytically.
@@ 2.2 BK equation for the first and second derivative of dipole amplitude
The BK equation is free of ultraviolet divergences, this follows from the probability conservation of the wave function where divergences of virtual diagrams cancel divergences of real diagrams when the emitted gluon $\bm{\bar{z}}_{\perp}$ is at the location of one of the two parents $\bm{\bar{x}}_{\perp}$ or $\bm{\bar{y}}_{\perp}$. Thus, the apparently singular behavior of the dipole kernel Eq. (2) when $\bm{\bar{z}}_{\perp}\rightarrow\bm{\bar{x}}_{\perp}$ or $\bm{\bar{z}}_{\perp}\rightarrow\bm{\bar{y}}_{\perp}$, is compensated by the combination of dipoles distributions $S_{\bm{\bar{x}}_{\perp}\bm{\bar{z}}_{\perp}}S_{\bm{\bar{z}}_{\perp}\bm{\bar{y}}_{\perp}}-S_{\bm{\bar{x}}_{\perp}\bm{\bar{y}}_{\perp}}]$ in the integrand of Eq. (1). As an illustration, for a Gaussian initial condition of the dipole distribution, the combination of dipoles $S_{\bm{\bar{x}}_{\perp}\bm{\bar{z}}_{\perp}}S_{\bm{\bar{z}}_{\perp}\bm{\bar{y}}_{\perp}}-S_{\bm{\bar{x}}_{\perp}\bm{\bar{y}}_{\perp}}]$ expanded around small separation $\bm{\bar{z}}_{\perp}-\bm{\bar{x}}_{\perp}$ scales as $\mathcal{O}\left((\bm{\bar{z}}_{\perp}-\bm{\bar{x}}_{\perp})^{2}\right)$, and the integrand in Eq. (1) is free of divergence at said point. This conclusion for UV-finiteness is unchanged when derivatives are taken w.r.t. the parent dipole size $\bm{\bar{r}}_{\perp}$, but due to the Leibniz rule individual terms may appear singular.
The regulator $\epsilon=5\cdot 10^{-4}\,\text{GeV}^{-1}$ is introduced in (4) to improve the numerical stability of the equations. We emphasize, that this regulator does not regulate the UV behavior of the equations, since they are already UV finite by themselves. In particular, $\epsilon$ can be smaller than the intrinsic UV scale provided by the lattice spacing. We checked that varying $\epsilon$ in a wide range from $5\cdot 10^{-5}\,\text{GeV}^{-1}$ to $5\cdot 10^{-3}\,\text{GeV}^{-1}$ had negligible impact on the results. On the other hand, $\epsilon$ stabilizes the evaluation of the derivatives of the dipole amplitude w.r.t. $r$ since no $0/0$ symbol is encounter in the limit $r_{z}\to r$, $\phi\to 0$. Note that the Jacobian for the polar coordinate provides $r_{z}$ that improves the behavior of the integrand when $r_{z}\rightarrow 0$. Hence, no regulator was needed there to evaluate the dipole amplitude and its first two derivatives w.r.t. $r$.
@@ 3.1 Forward-mode AD and dual numbers
In the present application, we employ a forward-mode implementation of AD because of its better performance for a small number of input parameters compared to reverse-mode [35, 36].
In the case of first derivatives, a real number is promoted to the abstract object called dual number, which is represented as an ordered pair of real numbers $(x,x^{\prime})$. The first component $x$ follows ordinary arithmetics, whereas rules for $x^{\prime}$ can be easily deduced from calculus on infinitesimal numbers. For example, sum, multiplication, $\sin$ function on dual numbers are given by:
$(x,x^{\prime})+(y,y^{\prime})$ $:=(x+y,x^{\prime}+y^{\prime}),$
$(x,x^{\prime})\cdot(y,y^{\prime})$ $:=(xy,xy^{\prime}+yx^{\prime}),$
$\sin(x,x^{\prime})$ $:=(\sin(x),x^{\prime}\cos(x)).$
Such rules can be defined for all elementary operations appearing in the algorithm. Second derivatives can be calculated in AD by defining dual numbers as triples of real numbers $(x,x^{\prime},x^{\prime\prime})$, where the arithmetic of the third component follows from the expansion of infinitesimal numbers up to the second order.
In our code, we replace real numbers with dual numbers defined as a C++ class as shown in Listing 1 in A. All necessary numeric primitives needed for the calculation are rewritten to operate on dual numbers and are contained in a single header file. We use operator overloading in C++ to handle dual numbers in order to increase the readability of the original source code, see Listing 2 in A for the example of a multiplication operator.